A mother is seven times as old as her daughter, in five years time she will be four times as old as her daughter. How old are they now?
Answer by mananth(16946)

A mother is seven times as old as her daughter, in five years time she will be four times as old as her daughter. How old are they now?
...
let daughter's age be x years
mother's age = 7x
five years later their ages will be
daughter =x+5
mother =7x+5
..
4(x+5)=7x+5
4x+20=7x+5
7x-4x=20-5
3x=15
/3
x=5 daughter's age
mother's age = 7*5 = 35 years
..
m.ananth@hotmail.ca